OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x answers 3x + 3x - 4 = 2 3x + 3 - 4x = 2 3x + 3x + 4 = 2

jhonyy9 (jhonyy9):

so than y+4=3x --- y=3x-4 and now substitute this value of y in first equation 3x +3x -4 =2

jhonyy9 (jhonyy9):

can you continue ?

jhonyy9 (jhonyy9):

do you need to solve this system ?

OpenStudy (anonymous):

no?

jhonyy9 (jhonyy9):

so 6x=6 --- x=6/6 --- x=1 y=3x-4 so y=3*1-4 so y=3-4 so y=-1

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

What work have you done?

jhonyy9 (jhonyy9):

ok so indifferent but your first answer is right sure ok ?

OpenStudy (anonymous):

ok